# BA-LoRA: Bias-Alleviating Low-Rank Adaptation to Mitigate Catastrophic Inheritance in Large Language Models

### v3

@@ 4.3.1 Analysis of the NLG and NLU Performance of BA-LoRA
BERT-L LoRA 87.24 93.19 90.10 64.73 93.13 90.94 73.14 90.63 85.39
BA-LoRA 89.72 94.85 92.23 65.49 95.48 91.72 75.77 91.71 87.12
GPT-2-XL LoRA 85.28 95.38 86.17 50.63 89.42 88.56 72.29 89.27 82.13
BA-LoRA 88.14 96.52 89.23 52.76 91.26 89.95 74.57 90.83 84.16
@@
BERT-L LoRA 85.19 93.49 89.93 63.49 92.32 87.73 73.65 90.57 84.55
BA-LoRA 87.91 94.18 90.62 65.81 93.04 89.06 75.41 91.21 85.91
GPT-2-XL LoRA 87.02 95.11 86.81 60.95 91.77 87.59 78.76 89.25 84.66
BA-LoRA 89.58 96.40 88.18 63.11 92.68 88.62 81.21 90.37 86.27
@@ 4.3.2 Analysis on mitigate noisy data
Figure 1: t-SNE Visualizations Comparing Last Hidden Layer Features of BERT-L and GPT-2-XL Fine-Tuned with LoRA and BA-LoRA on a Mini Subset of the GLUE Dataset.
@@ 4.3.3 Analysis on Mitigating Imbalanced Data
Figure 2: Ablation study results of BA-LoRA regularizations on the GSM8K and MATH datasets. Here, "$\mathrm{Reg}$" stands for "Regularization" and "$\mathit{w/o}\ \mathrm{Reg}$" means "without regularization". $\mathcal{L}_{\mathrm{CR\_NLG}}$, $\mathcal{L}_{\mathrm{DR\_NLG}}$, and $\mathcal{L}_{\mathrm{SVDR\_NLG}}$ denote the application of only the corresponding regularization, and "BA-LoRA" refers to the baseline using all regularizations.
Table 5: Ablation results of BA-LoRA regularization on NLU tasks. Here, "$\mathrm{Reg}$" stands for "regularization", and "$\mathit{w/o}\ \mathrm{Reg}$" means "no regularization". $\mathcal{L}_{\mathrm{CR\_NLU}}$, $\mathcal{L}_{\mathrm{DR\_NLU}}$, and $\mathcal{L}_{\mathrm{SVDR\_NLU}}$ mean that only the corresponding regularization is applied, and "BA-LoRA" refers to the baseline using all regularization. The best and second-best results are highlighted in bold and underline.
Method MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg
$\mathit{w/o}\ \mathrm{Reg}$ 90.47 95.81 91.48 72.27 94.41 92.21 87.14 91.93 89.47
$\mathcal{L}_{\mathrm{CR\_NLU}}$ 90.84 96.27 91.65 72.68 94.64 92.47 87.59 92.11 89.78
$\mathcal{L}_{\mathrm{DR\_NLU}}$ 90.77 96.09 91.81 72.44 94.44 92.41 87.37 91.89 89.65
$\mathcal{L}_{\mathrm{SVDR\_NLU}}$ 90.63 95.96 91.37 72.37 94.58 92.39 87.35 92.08 89.59
$BA-LoRA$ 90.92 96.25 91.83 72.79 94.84 92.59 87.87 92.15 89.91
4.3.4 Ablation Study
This ablation experiment aims to analyze the impact of three regularization terms ($\mathcal{L}_{\mathrm{CR\_NLG}}$, $\mathcal{L}_{\mathrm{DR\_NLG}}$, $\mathcal{L}_{\mathrm{SVDR\_NLG}}$) in BA-LoRA on model performance. The experiment selected three models, LLaMA-2-7B, Mistral-7B, and Gemma-7B, and evaluated them on GSM8K and MATH datasets. To ensure clear results, we only tested single regularization terms to reveal their independent contributions.
As shown in Figure 2, the model without regularization ("$\mathit{w/o}\ \mathrm{Reg}$") produced the lowest performance across both datasets. In contrast, introducing different regularization terms led to varying degrees of performance improvement. Specifically, $\mathcal{L}_{\mathrm{CR\_NLG}}$ contributes to the equilibrium gain for both datasets. The effect of $\mathcal{L}_{\mathrm{DR\_NLG}}$ on Gemma-7B is significantly stronger relative to $\mathcal{L}_{\text{CR\_NLG}}$. Ultimately, the model incorporating all three regularization terms achieved the highest performance on both datasets. These findings validate the regularization strategy and show that combining the terms in BA-LoRA further enhances model generalization. These findings confirm the effectiveness of the proposed regularization strategy, and further indicate that combining the regularization terms in BA-LoRA enhances the model’s generalization ability.
To further assess the impact of our proposed regularization terms on model performance, we conducted an ablation study focusing on three terms designed for NLU tasks in BA-LoRA: $\mathcal{L}_{\mathrm{CR\_NLU}}$, $\mathcal{L}_{\mathrm{DR\_NLU}}$, and $\mathcal{L}_{\mathrm{SVDR\_NLU}}$. We employed the DeBERTa-v3-base model and evaluated it across several tasks from the GLUE benchmark.
As shown in Table 5, the model without any regularization (denoted as "$\mathit{w/o}\ \mathrm{Reg}$") exhibited the lowest performance, achieving an average score of 89.47. Introducing each regularization term individually led to performance improvements across various tasks. Notably, $\mathcal{L}_{\mathrm{CR\_NLU}}$ consistently provided substantial gains, achieving an average score of 89.78, second only to the full BA-LoRA model. The term $\mathcal{L}_{\mathrm{DR\_NLU}}$ showed particularly strong performance on the MRPC task, while $\mathcal{L}_{\mathrm{SVDR\_NLU}}$ delivered balanced improvements across most tasks. Finally, BA-LoRA, incorporating all three regularization terms, achieved the highest overall performance with an average score of 89.91. These results demonstrate that each regularization term contributes to enhancing model performance, and their combination in BA-LoRA provides optimal generalization on NLU tasks.
@@ 5 Conclusion
This paper introduces Bias-Alleviating Low-rank Adaptation (BA-LoRA), a novel parameter-efficient fine-tuning method designed to mitigate catastrophic inheritance in pre-trained language models. BA-LoRA incorporates three key components: consistency regularization, diversity regularization, and singular value decomposition regularization. These components work in concert to preserve pre-training knowledge, enhance output diversity, and improve model generalization. Extensive experiments demonstrate that BA-LoRA consistently outperforms existing baselines on various NLG and NLU tasks while robust to noisy and imbalanced pre-training data. Furthermore, our ablation studies confirm the effectiveness of the three regularization terms both individually and in combination. These results highlight the potential of BA-LoRA as a general-purpose fine-tuning method for pre-trained language models and effectively address the key challenges of deploying these models in real applications.
@@ Appendix C More Experiments
Figure 4: Performance Comparison of LoRA and BA-LoRA Across Various Ranks.
@@ C.1 Analysis on Different Sizes and Types of Models
C.2 Evaluating the Performance of Different Ranks
We compare the performance of BA-LoRA, LoRA, and PiSSA at different ranks using LLaMA-2-7B and Mistral-7B-v0.1 models. Each method is fine-tuned for one epoch on the MetaMathQA-100K dataset with ranks ranging from 1 to 128 and evaluated on the GSM8K and MATH datasets. As Figure  4 shows, BA-LoRA consistently outperforms LoRA and PiSSA across all rank settings and datasets. As the rank increases, the performance of BA-LoRA and PiSSA surpasses full parameter fine-tuning. However, BA-LoRA performs better, especially on Mistral-7B-v0.1.
C.3 Impact of Regularization on LoRA Variants Performance
Table 10: Effect of regularization term in different LoRA variants. The best results are highlighted in bold.
@@ C.3 Impact of Regularization on LoRA Variants Performance
Full FT 6738M 49.05 7.22 21.34 35.59 4.91 23.62
LoRA 320M 42.47 5.60 17.03 31.48 4.62 20.24
LoRA + Reg 320M 49.26 6.84 20.16 32.75 4.75 22.75
DoRA 321M 42.12 6.28 16.97 22.07 4.53 18.46
PiSSA 320M 52.01 7.76 21.55 33.09 4.87 23.86
DoRA + Reg 321M 52.80 8.05 21.94 34.61 5.02 24.48
BA-LoRA 320M 53.83 9.13 23.58 36.86 5.11 25.70
@@
Figure 5: Impact of Different Regularization Hyperparameter Sizes for $\mathcal{L}_{\mathrm{task\_NLG}}$ on Fine-Tuned Model Performance, with Each Line Representing Results Using a Single Hyperparameter Value (Without Combinations).
C.4 Hyperparameter Analysis
This experiment aims to explore the impact of different sizes of regularization term hyperparameters on the performance of fine-tuned models for NLG tasks. The three regularization hyperparameters tested in the experiment are $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, with values selected from the range of $61e-$ 6, $51e-$ 5, $41e-$ 4, $31e-$ 3, and $21e-$ 2. We first fine-tune the LLaMA-2-7B model on the MetaMathQA dataset and then evaluate its performance on the GSM8K and MATH datasets.
As illustrated in Figure 5, the model achieves optimal performance on both datasets when $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to $1e^{-4}$. For smaller regularization values (e.g., $1e^{-6}$ or $1e^{-5}$), the regularization effect is minimal, leading to limited performance gains. In contrast, with larger regularization values (e.g., $1e^{-3}$ or $1e^{-2}$), the model’s performance decreases significantly. This decline occurs because higher regularization values impose stronger constraints on the training process, limiting the model’s capacity to learn and resulting in performance degradation.
C.5 t-SNE Visualizations of Feature Evolution during the Fine-tuning with LoRA and BA-LoRA
@@ C.4 Hyperparameter Analysis
Figure 6 shows that during LoRA fine-tuning of BERT-L, the feature separation is slow, with the class distributions remaining scattered and overlapping even towards the end. In contrast, Figure 7 demonstrates that with BA-LoRA fine-tuning, class separation begins earlier and is much clearer, ultimately forming a distinct "Y" shape with well-defined class boundaries.
@@ Appendix D More Discussions
D.1 Limitations
While BA-LoRA demonstrates significant improvements in mitigating catastrophic inheritance and enhancing model performance, several limitations warrant further investigation. Firstly, our evaluations primarily focus on English language tasks, which may limit the generalizability of our findings to other languages and specialized domains. Additionally, the computational overhead introduced by the consistency, diversity, and SVD regularizers adds complexity to the training process, potentially impacting efficiency. Furthermore, the impact of BA-LoRA on other forms of bias, such as fairness and societal stereotypes, remains unexplored. Lastly, the selection and weighting of regularization terms in BA-LoRA are fixed across different tasks, which may not be optimal for all scenarios.

### v4

@@ 4.4 Evaluating the Performance of Different Ranks
BERT-L LoRA 87.24±0.46 93.19±0.25 90.10±0.42 64.73±0.16 93.13±0.55 90.94±0.11 73.14±0.51 90.63±0.18 85.39
BA-LoRA 89.72±0.35 94.85±0.12 92.23±0.18 65.49±0.37 95.48±0.20 91.72±0.19 75.77±0.34 91.71±0.10 87.12
GPT-2-XL LoRA 85.28±0.39 95.38±0.46 86.17±0.31 50.63±0.54 89.42±0.06 88.56±0.24 72.29±0.36 89.27±0.16 82.13
BA-LoRA 88.14±0.28 96.52±0.38 89.23±0.19 52.76±0.26 91.26±0.11 89.95±0.37 74.57±0.22 90.83±0.07 84.16
@@
BERT-L LoRA 85.19±0.24 93.49±0.46 89.93±0.13 63.49±0.22 92.32±0.21 87.73±0.17 73.65±0.25 90.57±0.06 84.55
BA-LoRA 87.91±0.17 94.18±0.85 90.62±0.34 65.81±0.86 93.04±0.25 89.06±0.38 75.41±0.18 91.21±0.47 85.91
GPT-2-XL LoRA 87.02±0.05 95.11±0.44 86.81±0.11 60.95±0.38 91.77±0.04 87.59±0.07 78.76±0.03 89.25±0.13 84.66
BA-LoRA 89.58±0.08 96.40±0.13 88.18±0.04 63.11±0.14 92.68±0.08 88.62±0.19 81.21±0.48 90.37±0.27 86.27
@@ 4.4.1 Analysis on mitigate noisy data
Figure 2: t-SNE Visualizations Comparing Last Hidden Layer Features of BERT-L and GPT-2-XL Fine-Tuned with LoRA and BA-LoRA on the MNLI task of GLUE.
@@ 4.4.2 Analysis on Mitigating Imbalanced Data
Figure 3: Ablation Study of BA-LoRA Regularizations on GSM8K, MATH, and NLU Tasks. Here, “$\mathrm{Reg}$“ denotes “Regularization“, and “$\mathit{w/o}\ \mathrm{Reg}$“ indicates “without regularization“. $\mathcal{L}_{\mathrm{CR}}$, $\mathcal{L}_{\mathrm{DR}}$, and $\mathcal{L}_{\mathrm{SVDR}}$ represent the application of only the corresponding regularization, while “BA-LoRA“ refers to the baseline using all regularizations.
4.5 Ablation Study
This ablation study evaluates the impact of regularization terms in BA-LoRA on model performance for both NLG and NLU tasks, with results visualized in Figure 3. For NLG tasks, three models (LLaMA-2-7B, Mistral-7B, and Gemma-7B) were evaluated on the GSM8K and MATH using the regularization terms $\mathcal{L}_{\mathrm{CR\_NLG}}$, $\mathcal{L}_{\mathrm{DR\_NLG}}$, and $\mathcal{L}_{\mathrm{SVDR\_NLG}}$. The absence of regularization ($\mathit{w/o}\ \mathrm{Reg}$) consistently resulted in the lowest performance. Introducing individual terms led to varying improvements: $\mathcal{L}_{\mathrm{CR\_NLG}}$ provided balanced gains, $\mathcal{L}_{\mathrm{DR\_NLG}}$ showed stronger effects for Gemma-7B, and the combination of all three terms in BA-LoRA achieved the highest performance.
For NLU tasks, DeBERTa-v3-base was evaluated on the GLUE benchmark using the corresponding regularization terms $\mathcal{L}_{\mathrm{CR\_NLU}}$, $\mathcal{L}_{\mathrm{DR\_NLU}}$, and $\mathcal{L}_{\mathrm{SVDR\_NLU}}$. The reported results represent the average performance across eight tasks in the GLUE benchmark: MNLI, SST-2, MRPC, CoLA, QNLI, QQP, RTE, and STS-B. As shown in Figure 3, the model without regularization achieved the lowest average score of 89.47. Each of the three NLU-specific regularization terms consistently improved performance, with $\mathcal{L}_{\mathrm{CR\_NLU}}$ yielding a significant gain (average score of 89.78). The full BA-LoRA model, incorporating all three regularization terms, achieved the highest average score of 89.91, strongly demonstrating its effectiveness.
In summary, the ablation results demonstrate that the proposed regularization terms significantly enhance model performance for both NLG and NLU tasks. Each term contributes unique benefits, and their combination in BA-LoRA consistently achieves optimal results. This underscores the importance of integrating multiple regularization strategies to improve generalization and task-specific performance.
@@ 5 Conclusion
This paper introduces BA-LoRA, a novel parameter-efficient fine-tuning method designed to mitigate catastrophic inheritance in pre-trained language models. BA-LoRA incorporates three key components: consistency regularization, diversity regularization, and singular value decomposition regularization. These components work in concert to preserve pre-training knowledge, enhance output diversity, and improve model generalization. Extensive experiments demonstrate that BA-LoRA consistently outperforms existing baselines on various NLG and NLU tasks while robust to noisy and imbalanced pre-training data. Furthermore, our ablation studies confirm the effectiveness of the three regularization terms both individually and in combination. These results highlight BA-LoRA’s potential as a general-purpose fine-tuning method for pre-trained language models and effectively address the key challenges of deploying these models in real applications.
@@ C.1 Analysis on Different Sizes and Types of Models
C.2 Impact of Regularization on LoRA Variants Performance
Table 9: Effect of regularization term in different LoRA variants. The best and second-best results are highlighted in bold and underline.
@@ C.2 Impact of Regularization on LoRA Variants Performance
Full FT 6738M 49.13±0.21 7.29±0.22 21.20±0.30 35.59±0.25 4.91±0.01 23.62
LoRA 320M 42.47±0.29 5.60±0.35 17.03±0.61 32.77±0.46 4.62±0.11 20.50
LoRA + Reg 320M 51.91±0.30 8.63±0.38 21.11±0.03 33.85±0.16 4.75±0.35 23.83
DoRA 321M 42.12±0.16 6.28±0.19 16.97±0.08 22.07±0.38 4.53±0.47 18.41
PiSSA 320M 52.37±0.52 7.76±0.19 21.55±0.44 33.09±0.57 4.87±0.06 23.93
DoRA + Reg 321M 52.80±0.38 8.05±0.07 21.94±0.14 34.61±0.12 5.02±0.10 24.48
BA-LoRA 320M 54.53±0.41 9.21±17 23.58±0.25 36.86±0.31 5.11±0.05 25.86
@@
C.3 t-SNE Visualizations of Feature Evolution during the Fine-tuning with LoRA and BA-LoRA
@@
Figure 5 shows that during LoRA fine-tuning of BERT-L, the feature separation is slow, with the class distributions remaining scattered and overlapping even towards the end. In contrast, Figure 6 demonstrates that with BA-LoRA fine-tuning, class separation begins earlier and is much clearer, ultimately forming a distinct “Y” shape with well-defined class boundaries.
@@ Appendix D More Discussions
D.1 Computational Cost Analysis
Parameter-efficient fine-tuning (PEFT) methods are designed to minimize computational overhead while maintaining competitive performance relative to full fine-tuning. While PEFT methods are computationally efficient overall, specific components can introduce additional demands that warrant careful consideration. For instance, consistency regularization often necessitates an additional forward pass during training, leading to increased activation storage requirements. Similarly, operations like singular value decomposition (SVD) present scalability challenges for large matrices, as traditional SVD relies on eigenvalue computations that are computationally intensive and inherently CPU-bound. These factors highlight the importance of scalability, especially for high-dimensional models like LLaMA, where even relatively small matrices (e.g., $4096\times 4096$) demand substantial processing time.
To address these challenges, we introduce a randomized SVD approach, which achieves significant computational savings while preserving sufficient accuracy. Randomized SVD leverages random projections to approximate the dominant singular values and vectors, enabling faster computation and enhanced scalability for large matrices. This optimization ensures that computational demands remain manageable, even for models with high-dimensional weight matrices, thereby supporting scalability across diverse architectures.
To assess the practical implications of our approach, we conducted experiments measuring the average training time and memory usage across multiple runs. Specifically, we fine-tuned the DeBERTa-v3-base model on the full GLUE benchmark using an NVIDIA A40 GPU. The results for three PEFT methods—LoRA, PiSSA, and our proposed BA-LoRA—are summarized below:
  • LoRA: Averaged 14.45 hours of training time with an average memory usage of 6.3 GB.
  • PiSSA: Averaged 12.07 hours of training time with an average memory usage of 6.5 GB.
  • BA-LoRA: Averaged 15.82 hours of training time with an average memory usage of 7.1 GB.
Although BA-LoRA incurs marginally higher computational costs compared to LoRA and PiSSA, these increases stem primarily from its advanced regularization mechanisms, including the innovative use of randomized SVD. Crucially, these mechanisms effectively mitigate the catastrophic forgetting problem, resulting in substantial performance improvements. Considering the trade-off between computational overhead and model performance, the additional costs of BA-LoRA are both manageable and well-justified, especially in scenarios where achieving superior accuracy is paramount.